At market close on the date of your choosing, calculate the market cap (currency format to zero decimals with commas) of your company and two income (dividend paying) competitors. On the same date, calculate the annual dividend yield (percent format to two decimals) for the same three companies. Stock prices should be shown as currency format to two decimals. Provide one full page of original text comparing all 6 results and explaining which company would you invest in and why?
Answer #1

Solution :

Calculation of Market Capitalization and Dividend yield based on assumptive figures

Particulars Stock Price No. Of Shares Outstanding Divident Payout My Company $450.25 10,000,000 $50 A Company $380.15 15,000,

Explaination :

Comparing the market capitalization and dividend yield for all the three companies.

My Company

My company seems at mid cap level in comparison to other two companies. Although the market capitalization is higher then B Company, it has more potential to grow in comparison to A Company. There is lower risk in comparison to B Company. The dividend yield percentage of my company is almost near to the average of both A Company and B Company.

Divident yield percentage average between A Company and B Company = (Dividend Yield % of A Company + Divident Yield % of B Company)/2 = (15.78+7.68)/2 = 11.73

A Company

A Company seems at large cap in comparison to other two companies. Athough the market capitalization is higher then other two, there is low potential growth as compared to other two companies. The company is at highly volatile towards stock market.

B Company

B Company seems at small cap in comparison to other two companies. Although the stock price is higher then other two, there is high risk involved at low yield. The growth potential is very high in comparison to other two.

Considering all the above points, I personally would prefer a company wherein I can get good returns at moderate risk with potential to grow. My company is able to give average returns at moderate risk with a potential to grow in comparison to other two. So I would prefer keeping my investment in my company.